OpenStudy (anonymous):

What is the next step for 2cos(8θ/2)cos(-2θ/2) ???

OpenStudy (anonymous):

Divide 2 with 8 and -2 first..

OpenStudy (anonymous):

what do you want to prove ? :)

OpenStudy (anonymous):

I would like to solve 2cos(8θ/2)cos(-2θ/2) but i dont know how

OpenStudy (anonymous):

@waterineyes why do we have to divide by 2?

OpenStudy (anonymous):

"solve".... how ? ......the question ?!

OpenStudy (anonymous):

to simplify: \[2 \cos(\frac{8 \theta}{2}) \cos(\frac{-2 \theta}{2}) \implies 2 \cos(4 \theta) \cos(-\theta)\] \(cos( - \theta) = cos(\theta)\)

OpenStudy (anonymous):

how did the denominator disappear?

OpenStudy (anonymous):

Cancel 2 with 8..

OpenStudy (anonymous):

math is so hard.i hate being a student.

OpenStudy (anonymous):

Nooo! what do you want to do exactly ?@lsugano

OpenStudy (anonymous):

Ohi got the answer so its okay! its just that theres so many ways to find the answer and so many identities its making my calculations really hard! but i got it now!

OpenStudy (anonymous):

Im going to do a different question like this! If you dont mind, could you please check if imdoing it right?

OpenStudy (anonymous):

kk Sure !

OpenStudy (anonymous):

so far i got up to this!

OpenStudy (anonymous):

yes ! go on that's correct !

OpenStudy (anonymous):

:) I'm happy ! Go on !

OpenStudy (anonymous):

how do you get rid of the 2 in the denominator...? divide by 2?

OpenStudy (anonymous):

that's easy ! \[\frac{\frac{1}{2}}{2}\]=?

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

yes ! \[\frac{\frac{5\theta}{6}}{2}=??\]

OpenStudy (anonymous):

You're a great "student" :) !

OpenStudy (anonymous):

YAY! thank you!!

OpenStudy (anonymous):

you're welcome ! :) :)